Samples of a cast aluminum part are classified on the basis of surface finish (in microinches) and edge finish. The results of 101 parts are summarized as follows: edge finish excellent good surface excellent 81 finish !!goodーー118 10 Let A denote the event that a sample has excellent surface finish, and let B denote the event that a sample has excellent edge finish. If a part is selected at random, determine the following probabilities. Round your answers to three decimal places (e.g. 98.765) deimal places eg, espe (a) P(A) -To.83 (b) P(8)=1089 P(A)- (d) P(An8) (e) P(AUB)- P(A UB)
Answer #1

(a)

Edge finish excellent Edge finish good Total
Surface finish excellent 81 2 83
Surface finish good 8 10 18
Total 89 12 101

P(A) = P(Excellent surface finish) = 83/101 = 0.822

(b)

P(B) = P(Excellent edge finish) = 89/101 = 0.881

(c)

P(A') = P(Good surface finish) = 18/101 = 0.178

(d)

P(Aigcap B)= P(Excellent surface finish AND excellent edge finish) = 81/101=0.802

(e)

P(AUB)- P(A) +PUB) - PAB)

= 83/101 + 89/101 - 81/101 = 91/101= 0.901

(f)

P(A'igcup B)=P(A')+P(B)-P(A'igcap B)

= 18/101 + 89/101 - 8/101 = 99/101=0.980